# MTLoc: A Confidence-Based Source-Free Domain Adaptation Approach For Indoor Localization

### v1

@@ IV-C Ablation Studies
In this section, we observe the impact of hyper-parameter values, including the EMA’s alpha, $k$ of our $k$-NN-based correctional model, and the threshold we set for our $x$ and $y$-predicted pseudo labels. For this experiment, our MTLoc with confidence model is applied on the same datasets, testing $k$ values of 1, 2, and 3, along with alpha values of 0.6, 0.7, 0.75, and 0.8. Additionally, we applied $x$ and $y$ coefficients for thresholds of 3, 4, 5, and 4, 5, and 6, respectively. The validation and test losses were observed using 5-fold cross-validation. With an alpha of 0.8 and a $k$ value of 2 obtaining optimal results on the validation set for both the Cross and Square datasets using threshold coefficients of (5,4) and (4,2), we conducted an ablation study to assess the impact of $k$ and alpha, as shown in Tables II&III. Results indicate that $k=2$ provides the optimal number of neighbors, minimizing test errors for both $x$ and $y$ in both datasets. The experiment highlights that one neighbor is insufficient for correcting pseudo labels. Increasing $k$ from 1 to 2 or 3 does not significantly affect performance, confirming the reliability of both as hyperparameters. For EMA’s alpha, $\alpha=0.8$ achieves the best performance with more stability. In contrast, a higher $\alpha$ of 0.9 introduces over-smoothing, resulting in notable performance degradation. Additional details of the ablations on hyperparameters, including further experiments on threshold coefficients, are provided in Appendix B.
TABLE II: Ablation Results on $k$: The test MAE values and test distance errors (with standard deviations) are reported for the Cross and Square datasets at $K=1,2,3$, along with $\alpha=0.8$. For the Cross dataset, the $x$ and $y$ threshold coefficients are set to (5,4), respectively, while for the Square dataset, the thresholds are (4,2).
$\boldsymbol{k}$ Dataset Test MAE (x) $\downarrow$ Test MAE (y) $\downarrow$
1 Cross 1.96 $\pm$ 0.21 4.28 $\pm$ 0.16
Square 1.74 $\pm$ 0.14 4.35 $\pm$ 0.24
2 Cross 1.83 $\pm$ 0.12 4.17 $\pm$ 0.06
Square 1.70 $\pm$ 0.13 4.15 $\pm$ 0.05
3 Cross 1.88 $\pm$ 0.13 4.15 $\pm$ 0.04
Square 1.70 $\pm$ 0.09 4.19 $\pm$ 0.10
4 Cross 1.77 $\pm$ 0.12 4.73 $\pm$ 0.53
Square 1.68 $\pm$ 0.08 4.34 $\pm$ 0.15
@@
TABLE III: Ablation Results on $\alpha$: The test MAE values and validation distance errors are shown for the Cross and Square datasets at varying $\alpha$ values. The $x$ and $y$ threshold coefficients are set to (5,4) for the Cross and (4,2) for the Square datasets.
$\boldsymbol{\alpha}$ Dataset MAE (x) $\downarrow$ MAE (y) $\downarrow$
0.75 Cross 1.92 $\pm$ 0.23 4.44 $\pm$ 0.40
Square 1.73 $\pm$ 0.23 4.22 $\pm$ 0.16
0.80 Cross 1.83 $\pm$ 0.12 4.17 $\pm$ 0.06
Square 1.70 $\pm$ 0.13 4.15 $\pm$ 0.05
0.85 Cross 1.83 $\pm$ 0.26 4.47 $\pm$ 0.25
Square 1.89 $\pm$ 0.28 4.11 $\pm$ 0.14
0.90 Cross 3.17 $\pm$ 1.28 4.89 $\pm$ 0.87
Square 3.11 $\pm$ 0.27 5.80 $\pm$ 0.49
V Conclusions
@@
Data privacy, storage, transfer costs, and domain shifts are critical challenges in UDA, often affecting model performance in localization tasks. To address these concerns, an SFDA model was proposed for indoor localization that preserves data privacy and reduces transfer costs. In particular, the new MTLoc model is initialized by a localizer with the architecture of a 1D-CNN. MTLoc utilizes a student-teacher architecture that improves stability through an MT mechanism, where the teacher network generates pseudo labels. An enhanced certainty-based approach called ”MTLoc with confidence” selectively uses highly confident pseudo labels during training to improve performance. This technique refines the teacher network’s pseudo labels by including the two nearest confident labels. In our experimental validation, the MTLoc with confidence model provides lower localization errors compared to baseline methods, including adversarial UDA and SFDA models. The confidence-based approach notably reduced errors in the $y$ dimension and improved stability in both $x$ and $y$ with small variations, resulting in a more robust and reliable model.
Future works should explore semi-supervised techniques in environments with reference tags and validate them on additional datasets. Including labeled target data and fine-tuning models using reference tag information can significantly improve and optimize the performance.

### v2

@@ 5.3 Ablation Studies
In this section, we observe the impact of hyper-parameter values, including the EMA’s alpha, $k$ of our $k$-NN-based correctional model, and the threshold we set for our $x$ and $y$-predicted pseudo labels. For this experiment, our MTLoc with confidence model is applied on the same datasets, testing $k$ values of 1, 2, and 3, along with alpha values of 0.6, 0.7, 0.75, and 0.8. Additionally, we applied $x$ and $y$ coefficients for thresholds of 3, 4, 5, and 4, 5, and 6, respectively. The validation and test losses were observed using 5-fold cross-validation. With an alpha of 0.8 and a $k$ value of 2 obtaining optimal results on the validation set for both the Cross and Square datasets using threshold coefficients of (8,4) and (4,2), we conducted an ablation study to assess the impact of $k$ and alpha, as shown in Tables 2&3.
@@
Table 2: Ablation Results on $\alpha$: The test MAE values and validation distance errors are shown for the Cross and Square datasets at varying $\alpha$ values. The $x$ and $y$ threshold coefficients are set to (8,4) for the Cross and (4,2) for the Square datasets.
[…]
Table 3: Ablation Results on $\alpha$: The test MAE values and validation distance errors are shown for the Cross and Square datasets at varying $\alpha$ values. The $x$ and $y$ threshold coefficients are set to (8,4) for the Cross and (4,2) for the Square datasets.
[…]
5.4 Discussion
MTLoc is the first SFDA method developed for indoor localization, addressing the limited availability of source data under privacy or resource constraints. The approach outperforms state-of-the-art baselines, including UDA methods (DANN, ADDA, FiDO, and DAFI) and SHOT, a source-free classification method adapted for regression. MTLoc introduces several promising foundations for further extension. While the MTLoc model with confidence shows robust performance, it still falls short of an oracle model having access to labeled target data. A future direction could be studying the case where a small amount of labeled target samples is available during adaptation. In this setting, semi-supervised and weakly supervised domain adaptation techniques could be applied. In many real-world scenarios, it is feasible to collect a few labeled reference tags in the target domain. By including 5 to 10 labeled samples can significantly improve the model’s alignment with the target environment. These labeled points can serve as anchors to guide the model’s predictions toward the correct locations and help stabilize pseudo-labels. This, in turn, reduces the domain shift between source and target distributions. By using a small amount of supervision in an originally unsupervised SFDA setting, this extension could improve model accuracy while maintaining the low computational cost that makes MTLoc practical for large-scale deployment.
6 Conclusions
@@
Data privacy, storage, transfer costs, and domain shifts are critical challenges in UDA, often affecting model performance in localization tasks. To address these concerns, an SFDA model was proposed for indoor localization that preserves data privacy and reduces transfer costs. In particular, the new MTLoc model is initialized by a localizer with the architecture of a 1D-CNN. MTLoc utilizes a student-teacher architecture that improves stability through an MT mechanism, where the teacher network generates pseudo labels. An enhanced certainty-based approach called ”MTLoc with confidence” selectively uses highly confident pseudo labels during training to improve performance. This technique refines the teacher network’s pseudo labels by including the two nearest confident labels. In our experimental validation, the MTLoc with Confidence model achieves lower localization error than baseline methods, including DANN, ADDA, DAFI, FiDO, and the SFDA method SHOT. The confidence-based approach notably reduced errors in the $y$ dimension and improved stability in both $x$ and $y$ with small variations, resulting in a more robust and reliable model.
Acknowledgment
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), and Mitacs, with additional computational resources provided by the Digital Research Alliance of Canada.
7 Appendix A: Models Implementation
In this section, implementation details of the MTLoc and MTLoc with confidence, along with the three baselines, Adversarial UDA, source-only, and oracle models, are provided. In this study, a localizer, pre-trained on the INLAN’s Ceiling dataset, was designed and used in our source-only, UDA baseline and proposed SFDA approaches. An Adam optimizer with a learning rate of 0.001 was used for the optimization.
Source-only: In our source-only model, we used a localizer, having a feature extractor and a regressor.
Feature Extractor:
- Conv1D (filters = 64, kernel size = 2, activation = ‘ReLU’)
- Dropout (rate = 0.2)
- Conv1D (filters = 128, kernel size = 2, activation = ‘ReLU’)
- Dropout (rate = 0.2)
- Flatten
Regressor:
- Dense (units = 128, activation = ‘ReLU’)
- Dropout (rate = 0.2)
- Dense (units = 64, activation = ‘ReLU’)
- Dense (units = 2)
Adversarial UDA: In our Adversarial UDA baseline, we used a localizer having a feature extractor and regressor with the same structure as our source-only model, together with a discriminator.
Discriminator:
- Dense (units = 128, activation = ‘ReLU’)
- Dense (units = 64, activation = ‘ReLU’)
- Dense (units = 1, activation = ‘Sigmoid’)
MTloc: In our MTLoc model, our teacher and student networks have the same structure as our localizer. As the teacher network receives INLAN’s Cross and Squares datasets, the Student model obtains noisy versions of them as unlabeled target data with a variance of 0.1. By applying EMA with an alpha of 0.7, the teacher networks’ parameters aim to be close to the student’s parameters. In our confidence-based model, following the same methodology, a threshold was defined in which the average of the predicted two-dimensional positional coordinates $(x,y)$ is added to 5 and 3 times their respective standard deviation, correcting uncertain pseudo labels by applying kkNN (kk=2).
Oracle: In the Oracle model, we finetune our localizer, previously presented, on labeled target data (INLAN’s Cross and Square datasets).
8 Appendix B: Threshold Ablation Study
In this section, threshold hyperparameters of our MTLoc with confidence model are experimentally analyzed on both the Cross and Square datasets, to better understand the behavior of the proposed approach. Table A-4 presents the average test results of each experiment across 10 runs, along with the corresponding standard deviations obtained via 5-fold cross-validation. To define a confidence threshold for the positional coordinates, we independently scale the standard deviations of the $x$- and $y$-coordinate predictions with respect to their means and respective uncertainty distributions. This scaling factor was regarded as a tunable hyperparameter, with values of 4, 5, 6, and 8 evaluated for the $x$ positional coordinate, and 2, 3, and 4 for the $y$ positional coordinate. In our approach, using the cross and square datasets, we derived matrices of dimensions $1485\times 1$ and $1562\times 1$ for the $x$ and $y$ coordinates, respectively. Within the correction module, each training sample was processed 10 times, with random noise added in each iteration. This process resulted in matrices of dimensions $1485\times 10\times 2$ and $1562\times 10\times 2$, where the last dimension represents the $x$ and $y$ coordinates. After obtaining the outputs, we calculated the standard deviation across the 10 noise-augmented iterations for each sample, yielding $\sigma_{x}$ and $\sigma_{y}$ for the $x$ and $y$ coordinates, respectively. Our threshold is defined as:
$\textit{${T_{hr}}$}=\mu+c\cdot\sigma,$ (13)
where $\mu$ is the average of all the $\sigma_{x}$ or $\sigma_{y}$, $\sigma$ is the standard deviation of all $\sigma_{x}$ or $\sigma_{y}$, and $c$ is the scaling coefficient ($c_{x}$ or $c_{y}$). This threshold identifies samples with high uncertainty, enabling the correction module to modify these pseudo labels specifically. This approach ensures focus on improving the most unconfident predictions, thereby improving its overall robustness. Table IV shows an ablation study on threshold coefficients when $k$ and alpha are 2 and 0.8, respectively.
Table A-4: Performance of MTLoc model on two target datasets (Cross & Square) over threshold coefficients $c_{x}$ and $c_{y}$, for $k=2$ and $\alpha=0.8$.
[…]
9 Appendix C: Adapting SHOT To Regression
In this section, we discuss one of the baselines, SHOT [74], a classification-based SFDA method adapted to the regression setting, as shown in Fig. A-6. In the original setting, SHOT minimizes prediction entropy to ensure feature alignment. However, entropy is not applicable to regression tasks, as predictions are continuous positional coordinates. To address this, we maintain the original SHOT framework: a feature extractor $\mathcal{F}(\cdot;\theta_{s})$ and a frozen, source-pretrained regressor $\mathcal{G}(\cdot;\theta_{g})$. The model is trained to predict tag locations $y$ for unlabeled target data. During adaptation, only the feature extractor $\mathcal{F}(\cdot;\theta_{s})$ is trained, while $\mathcal{G}(\cdot;\theta_{g})$ remains frozen.
Figure A-6: Adaptation of SHOT for regression. The dashed black box indicates the teacher model, which can be optionally disregarded. Target inputs are augmented into two views: a weak view with small Gaussian noise $\mathcal{N}(0,\,0.01^{2})$ and a strong view with element-wise random masking ($p=0.10$) and Gaussian noise $\mathcal{N}(0,\,0.05^{2})$.
For each unlabeled target sample $z_{t}$, two augmented views are created: a weak view $x_{w}=a_{w}(z_{t})$ (small Gaussian noise, $\mathcal{N}(0,\,0.01^{2})$ and a strong view $x_{s}=a_{s}(z_{t})$ (element-wise random masking with probability $0.10$ having additive Gaussian noise, $\mathcal{N}(0,\,0.05^{2})$, with ${y}_{w}$ and ${y}_{s}$ as their predictions. To ensure prediction consistency across views, we minimize Eq. 14:
$\mathcal{L}_{\text{cons}}$ $=\mathbb{E}_{x\sim\mathcal{D}_{t}}\!\left[\left\|{y}_{w}-{y}_{s}\right\|_{2}^{2}\right].$ (14)
To increase the model’s stability during training, a second encoder $\mathcal{F}(\cdot;\theta_{t})$ is used as a teacher model, updated via EMA: $\theta_{t}\leftarrow\mathrm{EMA}(\theta_{s})$. In this student–teacher framework, the teacher encoder receives the weak, noisy data $x_{w}$ and produces target predictions. To ensure alignment between the student’s prediction on the strong view and the teacher’s outputs, we apply the teacher loss in Eq. 15:
$\mathcal{L}_{\text{teach}}$ $=\mathbb{E}_{x\sim\mathcal{D}_{t}}\!\left[\left\|\hat{y}_{s}-\hat{y}_{t}\right\|_{2}^{2}\right].$ (15)
Furthermore, as the source-domain prediction statistics and feature covariances are computed during pre-training, statistical alignment between source and target is ensured by matching the first- and second-order moments (mean and variance) of the student’s predictions across domains via the following prediction-statistics loss (Eq. 16):
$\mathcal{L}_{\text{pred-stat}}$ $=\left\|\mu_{t}-\mu_{s}\right\|_{2}^{2}+\left\|\sigma_{t}^{2}-\sigma_{s}^{2}\right\|_{2}^{2}.$ (16)
Additionally, to align feature-level covariance statistics across source features $\mathcal{F}_{s}$ and target features $\mathcal{F}_{t}$, we use the CORAL loss in Eq. 17 to align feature distributions across domains in this source-free setting:
$\mathcal{L}_{\text{coral}}$ $=\left\|\operatorname{Cov}(\mathcal{F}_{t})-\operatorname{Cov}(\mathcal{F}_{s})\right\|_{F}^{2},$ (17)
where $\|\cdot\|_{F}$ is the Frobenius norm. The total loss is defined as follows:
$\mathcal{L}_{\text{total}}=\lambda_{\text{cons}}\mathcal{L}_{\text{cons}}+\lambda_{\text{teach}}\mathcal{L}_{\text{teach}}+\lambda_{\text{stat}}\mathcal{L}_{\text{pred-stat}}+\lambda_{\text{coral}}\mathcal{L}_{\text{coral}}.$
This loss formulation balances consistency, statistical alignment, and teacher-guided supervision for effective SFDA. We use a learning rate of $10^{-3}$ and an exponential moving average coefficient of $0.995$. All loss weights are treated as hyperparameters and selected on a held-out validation split: $\lambda_{\text{cons}}=1.0$, $\lambda_{\text{teach}}=0.5$, $\lambda_{\text{stat}}=0.10$, and $\lambda_{\text{coral}}=0.02$.